# RAGAR: Retrieval Augment Personalized Image Generation Guided by Recommendation

### v1

@@ 5.1 Experimental Settings
We compare RAGAR with baselines through the following quantitative metrics: To evaluate personalization, we calculate the CLIP Personalization Score (CPS) and the CLIP Personalization Image Score (CPIS), which measure the similarity between the generated images and the text description or images representing user preferences, respectively. In addition, we calculate the LPIPS [?] and SSIM [?] to quantify the perceptual similarity. Furthermore, we compute the rank change $\mathbf{\Delta R}$ between the original and generated images. This metric emphasizes improvements for higher-ranked items by diminishing the weight for lower-ranked ones. A larger $\mathbf{\Delta R}$ indicates greater personalization in the image.
$\mathbf{\Delta R}=\frac{rk_{ori}-rk_{gen}}{1+rk_{ori}}$
where $rk$ denoted the rank determined by the ranking model. To evaluate semantic alignment, we calculate CLIP Score (CS), CLIP Image Score (CIS), LPIPS [?] and SSIM [?] for reference images.
@@
To make a fair comparison, all baselines are tuned with a fixed learning rate of $1e^{-5}$ and stable diffusion 1.5 is used as the image generator. For RAGAR, we set the learning rate at $1e^{-5}$. The number of retrieval items is fixed to 5. The number of noise is set at 3. All experiments are conducted on a single NVIDIA-A100 GPU.
5.2 Performance Comparison (RQ1)
The comparison between RAGAR and baselines is shown in Tab. 2. We can draw the following conclusions. First, traditional DM-based methods perform poorly on personalization metrics. Specifically, Glide and SD are the worst performers across five personalization metrics in all datasets. This is because both methods rely on the CLIP model to generate images, which struggles to capture deeper associations between texts. TI introduces a word embedding approach based on SD to capture user preferences. The stylized word embedding significantly improves personalization performance across all datasets (e.g., for $\Delta R$: -25.64 vs. -54.02, -2.58 vs. -3.20, -0.83 vs. -0.86). In terms of semantic alignment, all three methods perform well, and Glide achieves optimal results on SSIM across all datasets. This indicates that traditional methods tend to focus on reconstructing reference images.
Second, LLM-based methods demonstrate better performance, benefiting from their ability to understand and analyze text comprehensively. LaVIT, a multi-modal vision-language transformer, is designed to align textual and visual features effectively. Although it outperforms DM-based methods, it still falls short in personalization ($\Delta R$: -51.09 in the POG dataset). PMG, an LLM-based personalized image generation method, leverages LLMs to extract user preferences. This approach significantly enhances personalization performance, achieving optimal results compared to DM-based baselines. However, it heavily relies on the consistency with reference images, limited the personalization performance.
Third, our proposed method, RAGAR, achieves state-of-the-art (SOTA) performance due to the retrieval for sequence and the rank model for training across the three datasets. The performance boost is more pronounced on the ML dataset compared to POG and Sticker, likely because the amount of data in the front is significantly larger than that in the back. Compared to PMG, the second-best model, our method enhances consistency by semantically retrieving sequences and introducing evaluation metrics that better align with human judgment, resulting in stronger personalization performance.
@@ 5.2 Performance Comparison (RQ1)
Table 3: Human evaluation comparison on both personalization and semantic alignment.
5.3 Human Evaluation (RQ2)
To assess the personalization and semantics of RAGAR in real scenarios, we conduct human evaluation to compare it with PMG and original images. We invited 50 volunteers and set up two types of sorting tasks, with 50 cases for each dataset, covering personalization and semantics. For personalization, given 5 historical images along with the generated images, participants are asked to sort images in descending order of likelihood that the user would click them, based on historical click data. For semantics, given the reference image and generated images, participants are asked to sort them in descending order of how well they convey the same meaning as the reference image. As shown in Tab. 3, RAGAR outperforms the baselines on both metrics. This indicates RAGAR better reflects user preference while preserving the semantics.
Figure 4: Ablation study of the retrieval module and the rank rewards.
Figure 5: Examples of image generation across three datasets.
5.4 Ablation Study (RQ3)
We study the effectiveness of key components of RAGAR in the POG dataset, including retrieval and reflection module.
Effect of retrieval module.
@@ 5.4 Ablation Study (RQ3)
We assess the impact of retrieval module on improving semantic consistency and capturing user preferences by excluding it during training. Fig. 4(a) shows that: 1) noisy items in the historical sequence diminish the model’s ability to capture user preferences and interfere with semantics; 2) the retrieval module selects items relevant to user preferences, thereby improving the performance. Furthermore, we investigate the effect of varying retrieval number $k$ during training. The results reveal that $k=5$ yields the best results when historical sequence length is 20. At $k=0$ (no retrieval), performance is poor, while increasing $k$ initially improves performance but declines beyond $k=5$ due to reintroduced noise and redundancy.
Effect of rank rewards.
We investigate the role of rank rewards by excluding it while retaining the other loss functions during training, denoted as ‘w/o’. The results in Fig. 4(b) demonstrate that the model’s performance decreases without rank rewards. Adding the amount $r$ of sample noise $\epsilon$ boosts the model’s performance. Specifically, $r=5$ strikes the optimal balance between personalization and semantic alignment.
5.5 Case Study
We present cases of personalized images generated by RAGAR and PMG across three datasets: POG, ML-latest, and Sticker. Fig. 5 shows eight historical images, a reference image, and generated images for context. The orange dashed boxes highlight the relevant items retrieved by the RAGAR retrieval module, arranged together for clearer visualization. In clothing scenario, RAGAR generates red-and-blue patterned shoes, effectively capturing the user’s preference for shoes while filtering out irrelevant items, such as patent leather shoes, thereby outperforming PMG in focusing on user preferences. In the movie poster scenario, RAGAR reflects preferences for multiple characters appearing together and avoids irrelevant anime aesthetics, which PMG fails to exclude. In the sticker scenario, RAGAR maintains semantic coherence between a cat and a guitar while aligning with the user’s preference for pink ears and a brown cat style. It successfully filters out realistic style preferences, demonstrating its ability to balance personalization and semantics. Additional cases are in the supplementary material.
@@ 5.5 Case Study
Table 4: Quantitative comparison of the recommendation performances.
5.6 Auxiliary Generation(RQ4)
@@
Generating items beyond existing data can uncover user interests outside current data distribution [?]. Personalized images generated by RAGAR not only excel in display quality but also enhance recommendation tasks. We validate this using the multi-modal recommendation model MICRO[?]. We conduct experiments on two datasets with real interactions: POG and ML-latest. During training of rank model, reference images are replaced with those generated by RAGAR, PMG and original images ORI, respectively. The results in Tab. 4 reveal that generating images beyond original data distribution improves accuracy of the recommendation model. Both PMG and RAGAR outperform ORI across all metrics, demonstrating the benefit of generative methods. Notably, RAGAR shows a stronger capability in modeling user preferences compared to PMG, achieving significant improvements in metrics across both datasets, 6.8% (Recall@10) and 7.6% (NDCG@10) in POG, 2.67% (Recall@10) and 7.4% (NDCG@10) in ML-latest.
6 Conclusions and Future Work
In this paper, we proposed RAGAR for personalized image generation that balances personalization and semantics. With retrieval-augmented preferences, multi-modal alignment, and reflection-based optimization, RAGAR generates high-quality images that enhance both user experience and downstream recommendation performance. Extensive experiments demonstrate our method outperforms five baselines. For future work, we intend to unify preferences and noise in generation process to further enhance personalization.

### v2

@@ Methods
We compare RAGAR with baselines through the following quantitative metrics: To evaluate semantic alignment, we calculate CLIP Score (CS), CLIP Image Score (CIS), LPIPS (Zhang et al. 2018) and SSIM (Wang et al. 2004) for reference images. To evaluate personalization, we calculate the CLIP Personalization Score (CPS) and the CLIP Personalization Image Score (CPIS), which measure the similarity between the generated images and the text description or images representing user preferences, respectively. In addition, we calculate the LPIPS (Zhang et al. 2018) and SSIM (Wang et al. 2004) to quantify the perceptual similarity. Furthermore, we compute the rank change $\mathbf{\Delta R}$ between the original and generated images. This metric emphasizes improvements for higher-ranked items by diminishing the weight for lower-ranked ones. A larger $\mathbf{\Delta R}$ indicates greater personalization in the image.
$\mathbf{\Delta R}=\frac{rk_{ori}-rk_{gen}}{1+rk_{ori}}$ (16)
@@
To make a fair comparison, all baselines are tuned with a fixed learning rate of $1e^{-5}$ and Stable Diffusion 1.5 is used as the image generator. For RAGAR, we set the learning rate at $1e^{-5}$. The number of retrieval items is fixed to 5. The number of noise is set at 3. The hyperparameters for joint reflection are $\alpha=0.2,\beta=0.5,\gamma=0.3$. More training details can be found in the Appendix. All experiments are conducted on a single NVIDIA-A100 GPU.
Performance Comparison (RQ1)
Tab. 1 presents the comparison between RAGAR and baseline methods, from which several key observations can be made. First, traditional DM-based methods show poor performance on personalization metrics. Glide and SD rank lowest across all five personalization metrics on every dataset, largely because both rely on CLIP for image generation, which struggles to capture deeper semantic associations. TI, which augments SD with stylized word embeddings to model user preferences, significantly improves personalization (e.g., $\Delta R$: -25.64 vs. -54.02, -2.58 vs. -3.20, -0.83 vs. 0.86). In terms of semantic alignment, all three methods perform well, with Glide achieving the highest SSIM, indicating a tendency to reconstruct reference images.
Second, LLM-based methods perform better overall, owing to their stronger textual understanding. LaVIT, a multi-modal transformer, better aligns visual and textual features but still underperforms in personalization (e.g., $\Delta R$: -51.09 on POG). PMG, which extracts user preferences via LLMs, achieves better personalization than DM-based baselines but remains constrained by its dependence on reference image consistency.
Finally, RAGAR achieves state-of-the-art performance across all datasets, benefiting from sequence retrieval and a ranking-based training strategy. The gain is especially notable on the ML dataset, likely due to its larger volume of front-end user data. Compared to PMG, RAGAR further improves consistency through semantic retrieval and the use of evaluation metrics more aligned with human judgment, leading to superior personalization.
@@
Table 3: Human evaluation of RAGAR, PMG and origin images. RAGAR consistently outperforms both ORI and PMG in all metrics, achieving up to 29.46% improvement in personalization and 40.43% in semantic alignment.
Human Evaluation (RQ2)
To evaluate RAGAR’s effectiveness in personalization and semantic alignment, we conducted a human study comparing it with PMG and original images. Fifty participants completed two ranking tasks (50 cases per dataset): one on personalization—ranking generated images by their likelihood of being clicked based on 5 historical items—and one on semantics—ranking images by how well they convey the meaning of a reference image. As shown in Tab. 3, RAGAR consistently outperforms baselines on both tasks, demonstrating its superior ability to capture user preferences while preserving semantic content.
Ablation Study (RQ3)
@@
We study the effectiveness of key components of RAGAR in the POG dataset, including the retrieval and reflection module. Experiments on more datasets can be found in the Appendix. Effect of retrieval module. We assess the impact of the retrieval module on improving semantic consistency and capturing user preferences by excluding it during training. Fig. 4(a) shows that: 1) noisy items in the historical sequence diminish the model’s ability to capture user preferences and interfere with semantics; 2) the retrieval module selects items relevant to user preferences, thereby improving the performance. Furthermore, we investigate the effect of varying the retrieval number $k$ during training. The results reveal that $k=5$ yields the best results when the historical sequence length is 20. At $k=0$ (no retrieval), performance is poor, while increasing $k$ initially improves performance but declines beyond $k=5$ due to reintroduced noise and redundancy. Effect of rank rewards. We investigate the role of rank rewards by excluding it while retaining the other loss functions during training, denoted as w/o. The results in Fig. 4(b) demonstrate that the model’s performance decreases without rank rewards. Adding the amount $r$ of sample noise $\epsilon$ boosts the model’s performance. Specifically, $r=5$ strikes the optimal balance between personalization and semantic alignment.
Figure 4: The impact of the retrieval module and the rank reward mechanism. Performance improves as the number of retrieval items and noise samples increases, peaking at moderate values (around 5 for retrieval and noise), after which excessive inputs introduce noise and degrade results.
@@
Table 4: Application of personalized generation in recommendation domain. Compared to PMG and original images(ORI), RAGAR leads to an average improvement of approximately 7% in recommendation metrics.
Figure 5: Qualitative comparison of different methods. Three representative cases highlight how RAGAR incorporates retrieved features into the generation. Left: user interaction history (dashed boxes indicate retrieved items). Right: images generated by RAGAR and PMG.
Auxiliary Generation (RQ4)
@@
Generating items beyond the original data distribution can reveal user interests not captured by existing data (Li et al. 2024). RAGAR-generated personalized images not only offer high visual quality but also improve recommendation performance. We validate this using the multi-modal recommendation model MICRO (Zhang et al. 2023) on two real-world datasets: POG and ML-latest. In our experiments, we replace reference images with those generated by RAGAR, PMG, and original images (ORI) during rank model training. As shown in Tab. 4, both PMG and RAGAR outperform ORI across all metrics, highlighting the advantage of generative approaches. RAGAR achieves the best results, improving Recall@10 and NDCG@10 by 6.8% and 7.6% on POG, and by 2.67% and 7.4% on ML-latest, demonstrating its superior ability to model user preferences.
Case Study
We showcase personalized image generation results from RAGAR and PMG on the POG, ML-latest, and Sticker datasets (Fig. 5). Each example includes eight historical items, a reference image, and generated outputs, with orange dashed boxes marking items retrieved by RAGAR. In the clothing case, RAGAR produces red-and-blue shoes that match user preferences while filtering out irrelevant styles like patent leather, outperforming PMG. In the movie poster case, RAGAR reflects the user’s preference for multiple characters and avoids unrelated anime elements, which PMG fails to do. In the sticker case, RAGAR maintains semantic consistency (e.g., cat and guitar) and style (e.g., pink ears, brown fur), while excluding unwanted realistic styles. More examples are in the Appendix.
Related Works
Diffusion models (DMs)
(Rombach et al. 2021) have enabled high-quality image generation (Nichol et al. 2022; Xu et al. 2024b; Wei et al. 2024; Zeng et al. 2024; Wang et al. 2024; Zhang et al. 2025; He et al. 2025; Lu et al. 2025; Bi et al. 2024; Zhang et al. 2024) and video generation (Bi et al. 2025; Xu et al. 2025; Wang et al. 2025a; Cao et al. 2025; Feng et al. 2024; Shao et al. 2025), but personalization remains underexplored. Early methods (Gal et al. 2023; Ruiz et al. 2023) rely on text prompts, while others (Xu et al. 2024a; Yang et al. 2024; Shilova et al. 2023) combine user interactions and prompts to generate personalized content. Some approaches connect DMs with large language models (LLMs) (OpenAI et al. 2024; Touvron et al. 2023; Bai et al. 2023), such as (Koh, Fried, and Salakhutdinov 2023), and (Shen et al. 2024). However, (Shen et al. 2024) depends solely on consistency loss, often overfitting the reference image and ignoring user preferences. Multi-modal LLMs (Liu et al. 2023; Anil et al. 2023; Jin et al. 2024) generate personalized images from text but lack fine-grained preference modeling.
Recommendation systems
aim to deliver content aligned with user preferences, using retrieval (Borgeaud et al. 2022; Cui et al. 2024; Wang et al. 2025b) and ranking methods (Zhao et al. 2024; Geng et al. 2023; Zhang et al. 2023). Multi-modal approaches like (Cui et al. 2024) fuse text and image features for better reasoning, while (Geng et al. 2023) adds visual understanding. (Zhang et al. 2023) further models cross-modal user-item relationships. More detailed related work provide in the Appendix.
Conclusion
In the paper, we propose and validate two key assumptions in Sec. Introduction. Based on these insights, we introduce RAGAR, a novel framework that enhances personalization via retrieval-augmented preference and recommendation while preserving semantic consistency. We propose the Balance Calibrator to combine global and detailed preferences for guiding personalized generation. To address gradient propagation at the image level, we integrate recommendation signals and policy gradients into training. In future work, we aim to inject user preferences directly into the diffusion process to better align generation with user intent.